OpenStudy (anonymous):

choose the system of inequalities that match the graph below. A) 4x + y > 3 2x - 5y 15 B) 4x + y 3 2x - 5y > 15 C) 4x + y 3 2x - 5y < 15 D) 4x + y > 3 2x - 5y 15

OpenStudy (anonymous):

where is the graph

OpenStudy (dan815):

Have you found the equation of the dotted and thicker line

OpenStudy (anonymous):

i havent ... and I'm not sure how /: @dan815

OpenStudy (dan815):

Have you see this equation before y=mx+b or (y-y1)=m(x-x1)

OpenStudy (anonymous):

yes

OpenStudy (dan815):

Which one do you know how to use?

OpenStudy (anonymous):

umm y = mx + b

OpenStudy (dan815):

Okay so pick any 2 convenient points on the line. And find the slope for those 2 points and what the y-intercept must be(The y-intercepts are already given in fact).

OpenStudy (dan815):

The bottom line will be an less than or equal sigh, as its not dotted The above line, will be a greater only sign, since it is dotted and all the data is above the line.

OpenStudy (dan815):

y=mx+b,(m= solve, b= y-intercept)

OpenStudy (dan815):

Need any more clarification on how to find the slope of the line?